OpenStudy (nylicia_123):

Convert 0.0001 to a power of 10

OpenStudy (anonymous):

move the decimal to the right 10 times

OpenStudy (anonymous):

0001000000.

OpenStudy (danjs):

change it to a number between 1 and 10, not including 10, times 10^power

OpenStudy (danjs):

scientific notation

OpenStudy (nylicia_123):

thank u

OpenStudy (anonymous):

exactly what i said you are making it so difficult lol

OpenStudy (danjs):

the given decimal is the same as 10 to the -4 power

OpenStudy (danjs):

1*10^(-4)

OpenStudy (danjs):

also could call it 1 / 10000 1 ten thousandth = 0.0001 = 1/10000 = 10^-4

OpenStudy (nylicia_123):

okay so say i had 7.2 \[7.2\times 10 ^{-3}\] and i had to convert to standard form i wouls simply just ove the decimal to the left and my answer would be 7.002 @DanJS

OpenStudy (anonymous):

.0072

OpenStudy (anonymous):

you move it to the left three places

OpenStudy (anonymous):

if its positive more it to the right if it is negative move it to the left and depending on the number in the degree is how many places you move the decimal

OpenStudy (danjs):

recall, \[a^{-b} = \frac{ 1 }{ a^{+b} }\]

OpenStudy (danjs):

10^(-3) is the same as 1/10^3 or 1/1000

OpenStudy (danjs):

you are dividing by that power of ten, the number will become smaller by 4 decimal places

OpenStudy (danjs):

sorry, 3 decimal places, it is actually 10^-3 in the prob you gave

OpenStudy (danjs):

or just remember like Boots said left for negative powers, right for positive

OpenStudy (capnkookcx):

https://soundcloud.com/dann-gooas/record-0049

OpenStudy (clalgee):

Welcome to OpenStudy!